OpenStudy (anonymous):

6x - 8y + 6 = 36 6y = 4x - 24

OpenStudy (anonymous):

Firstly arrange the equations: 6x - 8y = 30 4x - 6y = 24

OpenStudy (anonymous):

Now you can do one thing more in this: \[3x - 4y = 15\] \[2x - 3y = 12\] Multiply first equation by 2 and second equation by 3 and then subtract them you will get y..

OpenStudy (anonymous):

thanks